OpenStudy (anonymous):

An aquarium is holding african chichilids is 5 inches higher than it is wide. Its lenght is 25 inches, and its volume is 7500 cubic inches. what are the height & the width of the aquarium

OpenStudy (pfenn1):

Volume = width x height x length if x is the width, then x+5 is the height V=x (x+5) 25=7500